# Navigating Challenges in Spatio-temporal Modelling of Antarctic Krill Abundance: Addressing Zero-inflated Data and Misaligned Covariates

### v1

@@ 4.1 Disaggregated Spatio-temporal Modelling
First, we fit the spatio-temporal Hurdle-Gamma model introduced in Section 3.1, including all covariates listed in Table 1 in the linear predictors for both the presence-absence and positive biomass components—i.e., Equations (1) and (2), respectively. Full details on this model are provided in Section SS2.1 (Supplementary Material). This approach was chosen to avoid the need for multiple model re-fits, as the associated computational cost was prohibitively high, despite the optimised inference specifications detailed in Section 3.1.2.
Table ST1 (Supplementary Material) shows the estimated coefficients, and Figure 4 shows the predicted values for both presence-absence and positive biomass components in 1997 and 2020. The corresponding results for the remaining years are shown in Figures SF2–SF21 (Supplementary Material). In the right-most plots of Figure 4, we masked out predicted values at locations with high uncertainty—specifically, where the standard deviation is greater than 3 (on the log scale, or approximately $20\text{g}/\text{m}^{2}$). As previously noted, the variability in the disaggregated data makes any spatio-temporal extrapolation beyond the observed locations extremely challenging using the modelling framework from Section 3.1. This also explains our focus on the region delineated by the green box (Figure 1). To address such a limitation and improve our ability to make predictions in non-observed areas, we shift to an aggregated analysis in Section 4.2.
Figure 4: Left column: observed acoustic krill biomass data. Middle column: estimated probability of non-zeros. Right column: predicted krill biomass ($\text{g}/\text{m}^{2}$), with predictions having a standard deviation greater than 3 (on the log scale) being masked out.
@@
However, before using these estimates to guide adjustments in data collection strategies, it is essential to assess their robustness under the Gaussianity assumption for the latent field. In Section SS2.1.1 (Supplementary Material), we conducted a sensitivity analysis by re-estimating the hyperparameters for observations generated from a latent non-Gaussian model. The results suggest that, while a potentially misspecified model may introduce a small bias in the range parameter (in particular, in our experiment, we noticed an upward bias of 5-10% for the parameter $\kappa$), the overall conclusions regarding acoustic survey sampling remain unaffected, as these differences are not substantial enough to meaningfully impact interpretation.
@@ 4.2 Aggregated Spatial Modelling
Following the approach of Warwick-Evans et al., (2022), who estimated krill abundance in the northern Antarctic Peninsula region, we fit a spatial Hurdle-Gamma model (as stated in Section 3.1.1) to an aggregated version of the original acoustic krill data. Specifically, we will aggregate the acoustic data both temporally and spatially into cells measuring $0.067^{\circ}$ longitude by $0.036^{\circ}$ latitude, corresponding to approximately $4\times 4$ km (as shown in the left column of Figure 6). Modelling this aggregated version of the data reduces issues of high spatial variability over short distances, making the Gaussianity assumption more reasonable and decreasing uncertainty in predictions beyond the observed area. While this approach sacrifices spatio-temporal resolution, it may offer valuable insights into the spatial distribution of krill biomass. In this section, we focus on making predictions within the red box (Figure 1), where there are more observations from KRILLBASE (net haul data), enabling us to compare and evaluate the accuracy of our extrapolated predictions—although, as discussed in Section 2.1, the net haul data do not measure krill biomass in the same way as the acoustic surveys and thus serve only as a proxy for the true spatial distribution.
@@
In this scenario, since the model is computationally much cheaper to fit, we can perform variable selection. Specifically, we will perform stepwise forward selection based on the Watanabe-Akaike Information Criterion (WAIC) (Watanabe,, 2013; Gelman et al.,, 2014). Additionally, we tested alternative models with simpler random effect structures (also using stepwise forward variable selection). However, the original model, i.e., the spatial Hurdle-Gamma model with linear predictors as in Equations (1) and (2), consistently outperformed these alternatives (Section SS2.2.1, Supplementary Material), reinforcing our choice to use it. Full details on the selected model are provided in Section SS2.2, where Table ST3 (Supplementary Material) shows the estimated coefficients. The covariates included in the linear predictors, in addition to the intercept, were as follows: for the presence-absence component, as in Equation (1), we considered chlorophyll, potential temperature, speed (satellite), surface temperature, mass flux, and residence time. For the positive krill biomass predictor, as in Equation (2), we considered depth, salinity, and surface temperature. Notably, some drifter-derived products improved the model’s performance, as indicated by the WAIC—suggesting that in situ and remotely sensed data may provide complementary information to the model.
@@
Figure 6 (middle column) shows the predicted krill biomass (log scale) based on the posterior mean. Model estimates were derived from the aggregated data (left column), with predictions based on covariate data from 2020—i.e., the upper limit of the acoustic krill biomass observations. The corresponding prediction uncertainty, represented by the $2.5^{\text{th}}$ and $97.5^{\text{th}}$ quantiles, is shown in Figure SF23 (Supplementary Material). The same figure shows the estimated probabilities (with uncertainty) of observing non-zero krill biomass.
@@
Figure 6: Left column: aggregated acoustic krill biomass data. Middle column: predicted krill biomass in 2020. Right column: KRILLBASE (net haul data), aggregated temporally and spatially. Krill biomass is in $\text{g}/\text{m}^{2}$.
@@ 5 Discussion
Finally, while our work demonstrates the potential of integrating multiple data sources, there are notable limitations. First, relying on acoustic surveys from a smaller region within Subarea 48.3 (green box in Figure 1) makes it challenging to extrapolate predictions beyond the observed window with reasonable uncertainty, particularly for the non-aggregated analysis (Section 4.1). Second, for the non-aggregated analysis, the assumption of Gaussianity for the underlying random effects in the Hurdle-Gamma model may lead to oversmoothed prediction maps. While this approach reduces computational burden during inference, exploring latent non-Gaussian models (e.g., Cabral et al., (2024)) in future work could offer greater flexibility and more effectively capture abrupt variations in the target ecological variable, even over short distances. Third, although we used net haul data (KRILLBASE) as a validation source for our extrapolated predictions, its spatial and temporal sparsity limits the reliability of conclusions about krill abundance at fine scales. Consequently, comparisons such as those presented in Figure 6 should be interpreted with caution, as noted in Section 4.2. Overall, additional extensions could further improve our model. An alternative approach is to explicitly account for spatially varying preferential sampling in the acoustic observations, combining this with data fusion techniques (Amaral et al.,, 2024; Zhong et al.,, 2024), as sampling locations are often influenced by ecological factors, potentially introducing bias into the analysis. Additionally, incorporating other in situ data sources, such as profiling floats (Roemmich et al.,, 2009), could provide complementary oceanographic measurements and refine the modelling of krill abundance.

### v2

@@ 4.1 Disaggregated Spatio-temporal Modelling
First, we fit the spatio-temporal Hurdle-Gamma model introduced in Section 3.1, retaining the complete random-effect structure and including all covariates listed in Table 1 in the linear predictors for both the presence-absence and positive biomass components—i.e., Equations (1) and (2), respectively. Full details on this model are given in Section SS3.1 (Supplementary Material). This approach was chosen to avoid the need for multiple model re-fits, as the associated computational cost was prohibitively high, despite the optimised inference specifications detailed in Section 3.1.2. Lastly, although the formulation in Section 3.1 is well-defined, note that practical identifiability issues between the two spatio-temporal random effects can arise when data are insufficiently informative or when priors are poorly specified.
In this setting, we focus on characterising the presence-absence component, whose linear predictor is defined in Equation (1). Furthermore, as previously mentioned, we examine key hyperparameters of the spatio-temporal random effects to gain deeper insight into the structure and design of the acoustic surveys’ sampling strategy.
Table ST3 (Supplementary Material) shows the estimated coefficients, and Figure 4 shows the predicted values for both presence-absence and positive biomass components in 1997 and 2020. The corresponding results for the remaining years are shown in Figures SF4–SF23 (Supplementary Material). In the right-most plots of Figure 4, we masked out predicted values at locations with high uncertainty—specifically, where the standard deviation is greater than 3 (on the log scale, or approximately $20\text{g}/\text{m}^{2}$). As previously noted, the variability in the disaggregated data makes any spatio-temporal extrapolation beyond the observed locations extremely challenging using the modelling framework from Section 3.1. This also explains our focus on the region delineated by the green box (Figure 1). To address such a limitation and improve our ability to make predictions in non-observed areas, we shift to an aggregated analysis in Section 4.2.
Figure 4: Left column: observed acoustic krill biomass data. Middle column: estimated probability of non-zeros. Right column: predicted krill biomass ($\text{g}/\text{m}^{2}$), with predictions having a standard deviation greater than 3 (on the log scale) being masked out. The two rightmost columns are based on the mean of the corresponding predictive distributions.
@@
However, before using these estimates to guide adjustments in data collection strategies, it is essential to assess their robustness under the Gaussianity assumption for the latent field. In Section SS3.1.1 (Supplementary Material), we conducted a sensitivity analysis by re-estimating the hyperparameters for observations generated from a latent non-Gaussian model. The results suggest that, while a potentially misspecified model may introduce a small bias in the range parameter (in particular, in our experiment, we noticed an upward bias of 5-10% for the parameter $\kappa$), the overall conclusions regarding acoustic survey sampling remain unaffected, as these differences are not substantial enough to meaningfully impact interpretation.
@@ 4.2 Aggregated Spatial Modelling
Following the approach of Warwick-Evans et al., (2022), who estimated krill abundance in the northern Antarctic Peninsula region, we fit a spatial Hurdle-Gamma model (as stated in Section 3.1.1) to an aggregated version of the original acoustic krill data. Specifically, we aggregate the acoustic data both temporally (all data from 1997 to 2020, Section 2.1.1) and spatially by taking the mean of observations within cells measuring $0.067^{\circ}$ longitude by $0.036^{\circ}$ latitude, corresponding to approximately $4\times 4$ km (as shown in the left column of Figure 6). Modelling this aggregated version of the data reduces issues of high spatial variability over short distances, making the Gaussianity assumption more reasonable and decreasing uncertainty in predictions beyond the observed area. While this approach sacrifices spatio-temporal resolution, it is explicitly aimed at providing a broad, large-scale picture of krill biomass across a wider area. Thus, in this section, we focus on making predictions within the red box (Figure 1), where there are more observations from KRILLBASE (net haul data), enabling us to compare and evaluate the accuracy of our extrapolated predictions—although, as discussed in Section 2.1, the KRILLBASE covers a different temporal range compared to the acoustic krill data (see Figure SF26, Supplementary Material). Additionally, the net haul data do not measure krill biomass in the same way as the acoustic surveys and thus serve only as a proxy for the true spatial distribution.
@@
In this scenario, since the model is computationally much cheaper to fit, we can perform variable selection. Specifically, we perform stepwise forward selection based on the Watanabe-Akaike Information Criterion (WAIC) (Watanabe,, 2013; Gelman et al.,, 2014). Additionally, we tested alternative models with simpler random effect structures (also using stepwise forward variable selection). However, the original model, i.e., the spatial Hurdle-Gamma model with linear predictors as in Equations (1) and (2), consistently outperformed these alternatives (Section SS3.2.1, Supplementary Material), reinforcing our choice to use it. Full details on the selected model are provided in Section SS3.2, where Table ST5 (Supplementary Material) shows the estimated coefficients. The covariates included in the linear predictors, in addition to the intercept, were as follows: for the presence-absence component, as in Equation (1), we considered chlorophyll, potential temperature, speed (satellite), surface temperature, mass flux, and residence time. For the positive krill biomass predictor, as in Equation (2), we considered depth, salinity, and surface temperature. Notably, some drifter-derived products improved the model’s performance, as indicated by the WAIC—suggesting that in situ and remotely sensed data may provide complementary information to the model.
@@
Figure 6 (middle column) shows the predicted krill biomass (log scale) based on the posterior mean. Model estimates were derived from the aggregated data (left column), with predictions based on covariate data from 2020—i.e., the most recent year for which acoustic krill biomass observations are available. The corresponding prediction uncertainty, represented by the $2.5^{\text{th}}$ and $97.5^{\text{th}}$ quantiles, is shown in Figure SF25 (Supplementary Material). The same figure shows the estimated probabilities (with uncertainty) of observing non-zero krill biomass.
@@
Figure 6: Left column: aggregated acoustic krill biomass data. Middle column: predicted krill biomass in 2020 (based on the mean of the predictive distribution). Right column: KRILLBASE (net haul data), aggregated temporally and spatially. Krill biomass is in $\text{g}/\text{m}^{2}$.
@@ 5 Discussion
Finally, while our work demonstrates the potential of integrating multiple data sources, there are notable limitations. First, relying on acoustic surveys from a smaller region within Subarea 48.3 (green box in Figure 1) makes it challenging to extrapolate predictions beyond the observed window with reasonable uncertainty, particularly for the non-aggregated analysis (Section 4.1). Second, for the non-aggregated analysis, the assumption of Gaussianity for the underlying random effects in the Hurdle-Gamma model may lead to oversmoothed prediction maps. While this approach reduces computational burden during inference, exploring latent non-Gaussian models (e.g., Cabral et al., (2024)) in future work could offer greater flexibility and more effectively capture abrupt variations in the target ecological variable, even over short distances. Third, in Section 3, we model each covariate with a single linear term; introducing a compact non-linear basis would add flexibility without otherwise modifying the hierarchical framework. Fourth, although we used net haul data (KRILLBASE) as a validation source for our extrapolated predictions, its spatial and temporal sparsity limits the reliability of conclusions about krill abundance at fine scales. Consequently, comparisons such as those presented in Figure 6 should be interpreted with caution, as noted in Section 4.2. Overall, additional extensions could further improve our model. Given that vessel routes are generally predefined based on prior ecological knowledge—and that sampling locations are determined along these routes—methods that explicitly model the resulting bias, such as that of Amaral et al., (2024), can be combined with data fusion techniques to improve inference; Zhong et al., (2024) provides one such approach. Additionally, incorporating other in situ data sources, such as profiling floats (Roemmich et al.,, 2009), could provide complementary oceanographic measurements and refine the modelling of krill abundance.